OpenStudy (anonymous):

Is the diff equation linear? Separable? Both? Neither? X dx/dt + t^2 x = sin t I have it to be separable but not linear??

OpenStudy (unklerhaukus):

\[x \frac{\text dx}{\text dt} + t^2 x = \sin t\] \[\frac{\text dx}{\text dt} + t^2 = \frac{\sin t}x\] \[\frac{\text dx}{\text dt} = \frac{\sin t}x-t^2\]

OpenStudy (unklerhaukus):

how did you separate the variables?